# Concept-1K: A Novel Benchmark for Instance Incremental Learning

### v2

@@ 2 Instance Incremental Learning
Figure 2: The analysis of memorization (top row) and generalization (bottom row) accuracy on Concept-1K. The backbone model is in {Pythia-70M, 160M, 410M, 1B, 1.4B, 2.8B}. The pretraining step is in {0, 16, 128, 1000, 10000, 143000 (final version)}. Each point represents the result of sequential finetuning under the IIL scenario. The detailed results with standard deveriation are provided in Table 7, 8, 9, 10, 11, 12, 13, and 14.
Concept-1K To avoid the data leakage issue, we collect novel concepts from six domains, including economy, culture, science and technology, environment, education, and health and medical. To begin with, we generate 600 concepts for each domain with GPT4. We then manually filter out outdated, vague, or imaginary concepts and select the latest, most specific, and most informative concepts, resulting in a total of 1023 concepts. The concept list is provided in Table 17. Next, we use triplets to represent “knowledge” and prompted GPT4 to construct 20 triplets for each concept with the relation in ConceptNet Speer et al. (2017). To avoid conflict, we filter out the triplets with the same concept and relation. In addition, we filter out the triplets with relations such as “RelatedTo” and “HasContext” to ensure specificity. Finally, we utilize GPT4 to convert each triplet into a pair of training and test instances in QA format. We provide the examples and statistics of Concept-1K in Table 1, 15, and Figure 5, 6.
3 Experimental Results
We divide 1023 concepts of Concept-1K into 10 groups for IIL. We explore the impact of the model scale, pretraining, buffer size, and finetuning methods (FULL or LoRA) on the performance under the IIL scenario. We also investigate whether in-context learning is qualified to adapt PLMs to novel concepts in Concept-1K. In addition, we analyze the IIL performance at the concept level in the Appendix C.4. The detailed experimental settings, implementation details, and introduction of baselines are provided in the Appendix B.
Table 2: The forgetting of PLMs with different scales. The pretraining step is final and buffer size is 0.
@@ 3 Experimental Results
The Role of Model Scale and Buffer Size. The results are summarized in Figure 2 and Table 2. The findings are summarized as follows: (1) Figure 2 (a) and (e) show that a larger buffer size or a larger PLM improve the accuracy of memorization and generalization. (2) The model scale determines the upper limit of the generalization performance. (3) However, the memorization accuracy of the 2.8B model is still unsatisfactory when no buffer is used. It indicates that the billion-parameter PLMs still suffer from catastrophic forgetting, and data replay is an effective technique for alleviating it. (4) Table 2 shows that when PLMs become larger, the memorization forgetting decreases while the generalization forgetting increases. It indicates that larger PLMs forget less training samples while forget more generalized knowledge since more knowledge are generalized.
@@
The Role of Pretraining. Figure 2 (b) shows that the memorization performance increases at the early stage of pretraining (step 0 - step 10000). It indicates that pretraining enhances the memorization ability of PLMs to novel concepts. However, when performing more pretraining steps, the memorization performance will be degraded. In contrast, Figure 2 (f) shows that the generalization performance increases monotonically for PLMs larger than 160M. The reason may be that PLMs will gradually learn to extract the knowledge behind the text during the pretraining process instead of simply remembering specific texts. In addition, especially for larger models, pretraining is more beneficial to generalization ability rather than memorization ability. Furthermore, Figure 2 (c)(d)(g)(f) indicates that a larger buffer size improves both memorization and generalization ability for all pretraining steps. The improvement of memorization ability is more significant than that of generalization ability.
Method MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$)
SEQ 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16
EWC Kirkpatrick et al. (2017) 59.83±0.62 18.09±0.28 62.46±0.69 15.07±0.33
LAMOL_g Sun et al. (2020) 58.76±0.53 15.35±0.46 64.64±0.59 13.61±0.47
LAMOL_t Sun et al. (2020) 58.29±2.17 15.24±0.37 66.66±2.42 14.05±0.38
L2KD Chuang et al. (2020) 28.34±0.29 10.87±0.01 32.45±0.63 8.55±0.42
PCLL Zhao et al. (2022) 61.94±1.41 20.06±0.42 63.04±1.57 16.27±0.37
LFPT5 Qin and Joty (2022) 0.63±0.04 0.84±0.01 0.04±0.03 0.03±0.01
LAMOL_KD Zheng et al. (2023) 72.33±0.45 18.20±0.37 49.25±0.32 10.61±0.42
REPLAY (buffer size=2000) 77.31±0.22 22.48±0.29 46.97±1.50 10.57±0.24
REPLAY (buffer size=Alll) 99.01±0.11 25.70±0.44 0.70±0.16 1.44±0.88
Performance of SOTA Methods. We compare 8 rehearsal-free methods and data replay (REPLAY) under the IIL scenario. Table 3 indicates that although existing methods have improved sequential finetuning (SEQ), there is still a gap in the data replay with only 2000 samples (about 12 % of the total sample). The gap in memorization accuracy is much greater than generalization accuracy, indicating that PLMS still suffers from catastrophic forgetting under the challenging IIL scenario. The result urges us to design a more powerful IL algorithm to reduce the need for data replay.
Table 4: The IIL performance of full finetuning (FULL) and LoRA on various backbones.
MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$)
Pythia-410M (FULL) 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16
Pythia-410M (LoRA) 15.72±0.72 6.58±0.15 30.97±1.15 2.96±0.14
Pythia-2.8B (LoRA) 36.56±0.93 10.93±0.17 83.94±0.64 6.04±0.23
Vicuna 7B (LoRA) 42.28±0.25 16.74±0.33 78.32±0.54 7.69±0.47
LLaMa 7B (LoRA) 41.76±0.27 16.20±0.18 80.55±0.39 8.08±0.31
LLaMa 13B (LoRA) 48.90±0.68 22.67±0.30 74.55±0.46 12.26±0.18
LoRA v.s. Full Finetuning. LoRA Hu et al. (2021) only finetunes a small proportion of parameters (usually less than 1%) and effectively adapt PLMs to specific domains. We compare the performance of full finetuning with LoRA in the IIL setting. Table 4 shows that LoRA limits the ability to memorize new knowledge or generalize to new instances. More specifically, the memorization and generalization accuracy of Pythia-410M is higher than LLaMa 7B (LoRA). This reminds us that when we need PLM to learn much new knowledge, we should prioritise using full fine-tuning instead of LoRA.
Table 5: The accuracy of in-context learning on the full Concept-1K (not IIL). “Rand.”, “Same Conc.”, and “Same Know.” represent that the demonstration samples are selected randomly, or from the instances related to the same concept, or from the instance related to the same knowledge (i.e., the same training-test pair).
1 Shot 5 Shot
Rand. Same Conc. Same Know. Rand. Same Conc. Same Know.
GPT 3.5 6.60 8.20 51.60 8.60 13.00 74.80
GPT 4 10.20 10.40 76.60 7.40 21.80 86.20
In-Context Learning. Due to computing resources and privacy concerns, it is impractical to fully or partially fine-tune SOTA PLMs such as GPT4. We evaluate the in-context learning performance on the whole Concept-1K to investigate the ability to learn novel concepts. Table 5 indicates that the performance is unsatisfactory for all PLMs when the demonstration instances are less related to the test instance. Additionally, we can improve the performance by leveraging demonstration instances that share the same knowledge as the test instances. In other words, PLMs achieves superior performance only when the demonstration instances containing exactly the same knowledge as the test sample. Therefore, in-context learning can not meet the goal of adapting PLMs to new knowledge.
@@ 4 Conclusion
Incremental learning is a research field that has developed rapidly in recent years. However, the existing IL scenario and datasets are not suitable for SOTA PLMs and even give an illusion that PLMs do not suffer from catastrophic forgetting. This study presents a novel IL scenario called instance-incremental learning and a novel dataset called Concept-1K, supporting an order of magnitude larger steps for IL. We hope our study will further inspire follow-up works to explore the catastrophic forgetting in PLMs and to design more powerful techniques for finetuning and incremental learning.
@@
There are two limitations of this research: (1) The knowledge of Concept-1K is defined in the form of triplets, which can not cover the knowledge in a broad sense. (2) Apart from the experiments of in-context learning, other experiments are conducted on PLMs with less than 13B parameters. The conclusion of these experiments may not hold when finetuning SOTA PLMs such as GPT4 with more than 100B parameters.
@@
Appendix A Evaluation Metrics
The memorization accuracy (MA) is defined as follows:
$MA=\frac{1}{T}\sum_{t=1}^{T}\mathcal{A}_{t},$ (1)
where $T$ is the number of incremental steps. $\mathcal{A}_{t}$ represents average accuracy on the training instances from all learned concepts. Specifically, $\mathcal{A}_{t}$ is computed as follows:
$\mathcal{A}_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i},$ (2)
where $a_{t,i}$ represents the accuracy evaluated on the task $i$ after training the model incrementally from concepts belonging to group $1,\cdots,t$. The accuracy is calculated as the exact match between the model output and the target output.
The memorization forgetting (MF) is computed as the average accuracy on all training instances of all learned concepts:
$MF=\frac{1}{T-1}\Sigma_{i=1}^{T-1}[max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}].$ (3)
Concretely, $max_{j<T}(\{a_{j,i}\}_{j})$ represents the highest accuracy of task $i$ since it has been learned. $a_{T,i}$ represents the accuracy of task $i$ at step $T$. $[max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}]$ computes the decrease of the accuracy of task $i$ at incremental step $T$.
The generalization accuracy (GA) and generalization forgetting (GF) are computed in a similar way except that the model is evaluated on the test set instead of the training set.

### v3

@@ 3.4 RQ4: What is the most effective and efficient method for IL of LLMs?
Figure 4: The analysis of memorization (top row) and generalization (bottom row) accuracy on Concept-1K. The backbone model is in {Pythia-70M, 160M, 410M, 1B, 1.4B, 2.8B}. The pretraining step is in {0, 16, 128, 1000, 10000, 143000 (final version)}. Each point represents the result of IL. The detailed results with standard deveriation are provided in Table 13, 14, 15, 16, 17, 18, 19, and 20.
Table 8: The forgetting of LLMs with different scales. The pretraining step is final and buffer size is 0.
@@
3.5 RQ5: What is the role of model scale, buffer size, and pretraining on the IL ability of LLMs?
Main Findings 5: Larger model scale, buffer size, and more pretraining steps all lead to better IL ability.
Given that data replay is effective, we next explore how model scale, buffer size, and pretraining influence the incremental learning ability of LLMs. The results are summarized in Figure 4 and Table 8. Detailed results corresponding to Figure 4 can be found in Tables 13, 14, 15, 16, 17, 18, 19, and 20.
@@ 3.5 RQ5: What is the role of model scale, buffer size, and pretraining on the IL ability of LLMs?
Model Scale. The model scale determines the upper limit of generalization performance. Table 8 shows that as LLMs become larger, the memorization forgetting decreases while the generalization forgetting increases. This indicates that larger LLMs forget fewer training samples but more generalized knowledge, as they generalize more knowledge.
Buffer Size. Figures 4 (a) and (e) show that a larger buffer size or a larger LLM improves the accuracy of both memorization and generalization. However, the memorization accuracy of the 2.8B model remains unsatisfactory without a buffer. This suggests that even billion-parameter LLMs suffer from catastrophic forgetting, and data replay is an effective technique for mitigating it. Furthermore, Figures 4 (b)-(d), (f)-(h) indicate that a larger buffer size improves both memorization and generalization abilities across all pretraining steps, with the improvement in memorization ability being more significant than that in generalization ability.
@@
Pretraining. Figure 4 (b) demonstrates that memorization performance increases during the early stages of pretraining (step 0 - step 10000), indicating that pretraining enhances the memorization ability of LLMs for novel concepts. However, with more pretraining steps, memorization performance degrades. In contrast, Figure 4 (f) shows that generalization performance increases monotonically for LLMs larger than 160M. This may be because LLMs gradually learn to extract underlying knowledge from the text during pretraining, rather than merely remembering specific texts. Additionally, especially for larger models, pretraining is more beneficial to generalization ability than memorization ability.
(a) Sorted by MA
(b) Sorted by GA
Figure 5: The memorization accuracy and generalization accuracy of different concepts in Concept-1K. The concepts are sorted according to (a) memorization accuracy and (b) generalization accuracy respectively.
Table 9: The concepts with highest and lowest generalization accuracy.
Top 5 Bottom 5
Concept GA (%) MA (%) Concept GA (%) MA (%)
Peer-to-Peer Lending 82.47 100.00 Smart City Technologies 0.00 25.54
Letters of Credit 77.61 100.00 Orbital Mechanics 0.00 44.13
Streaming Services 74.43 100.00 Virtual Fitting Rooms 0.00 32.86
Carbon Neutral 71.04 85.71 Mars Missions 0.00 5.87
Interest Rate Hikes 70.82 100.00 Flexible Displays 0.00 26.38
3.6 RQ6: Are concepts learned equally?
Main Findings 6: Concepts that are well-defined and concrete are easier to memorized and generalized.
Finally, we explore whether LLMs learn all concepts equally. Are some concepts easier to learn? We analyze the memorization and generalization accuracy of various concepts in the Concept-1K dataset, using the LLaMa-2-7B model as the backbone. To mitigate the impact of task order, we aggregate the performance of the concepts in the fifth task after training on all tasks from 10 different task orders.
Figure 5 reveals a positive correlation between memorization accuracy and generalization accuracy, indicating that concepts easier to memorize are also easier to generalize, and vice versa. Our findings align with those of Toneva et al. (2018), which suggest that certain examples are unforgettable and their knowledge can be better generalized across datasets.
Table 9 highlights that concepts with the highest memorization accuracy tend to be well-defined and concrete, often related to established financial or technological terms. In contrast, concepts with the lowest memorization accuracy are often more complex, abstract, or emerging fields, which may explain the challenges in both memorization and generalization. This disparity underscores the importance of the nature of the concepts being learned and the inherent difficulty associated with them. Our findings are consistent with those of Toneva et al. (2018), which reveal that unforgettable images are easily recognizable, while the most forgotten examples exhibit more ambiguous characteristics.
(a) After Task 1 (Train)
(b) After Task 10 (Train)
(c) After Task 1 (Test)
(d) After Task 10 (Test)
Figure 6: The visualization of the memorized and generalized knowledge related to the “Brain-Computer Interface” in IL. The center node represents a concept, while the linked and unlinked edges indicate whether the corresponding training and test samples are answered correctly. The full results are in Figure 7 and 8.
Figure 6 visualize the memorized and generalized knowledge of one individual concept “Brain-Computer Interface” and illustrate the forgetting on knowledge during IL. The full results and further discussion are provided in Appendix H.
4 Related Work
We categorize existing studies on understanding the incremental learning ability of LLMs into three parts: (1) Understanding Forgetting, (2) Understanding Memorization, and (3) Applications in NLP. Due to space limitations, the detailed discussion is provided in the Appendix A.
Understanding Forgetting. Earlier studies, such as French (1999), assess catastrophic forgetting by measuring performance degradation on old tasks. Recently, studies Tao et al. (2023); Zheng et al. (2023b) use probing techniques to measure forgetting in incremental learning. Zheng et al. (2023b) uses probing techniques to show that LLMs have superior performance on evaluated datasets even before IL. Our work is inspired by Zheng et al. (2023b) and proposes a novel dataset to minimize the influence of data leakage issues.
@@ 5 Conclusion
In this paper, we introduce Concept-1K, a novel dataset designed to evaluate the IL capabilities of LLMs. Our comprehensive experiments reveal that LLMs still suffer from catastrophic forgetting and that LoRA, despite fine-tuning fewer parameters, limits the ability to learn and generalize new knowledge. We demonstrate that data replay is the most effective method for mitigating forgetting and highlight the significant roles of model scale, buffer size, and pretraining. These findings provide valuable insights into the strengths and limitations of LLMs in IL scenarios, offering a robust benchmark for future research.
@@
There are two limitations of this research: (1) The knowledge of Concept-1K is defined in the form of triplets, which can not cover the knowledge in a broad sense. (2) Apart from the experiments of in-context learning, other experiments are conducted on LLMs with less than 13B parameters. The conclusion of these experiments may not hold when finetuning SOTA LLMs such as GPT4 with more than 100B parameters.
@@
Appendix A Related Work
We categorize existing studies on understanding the incremental learning ability of LLMs into three parts: (1) Understanding Forgetting, (2) Understanding Memorization, and (3) Applications in NLP.
A.1 Understanding Forgetting
Earlier studies, such as French (1999); Kirkpatrick et al. (2017), assess catastrophic forgetting by measuring performance degradation on old tasks. Recently, studies Davari et al. (2022); Wu et al. (2021); Chen et al. (2023); Tao et al. (2023); Zheng et al. (2023b) use probing techniques to measure forgetting in incremental learning. For example, Davari et al. (2022) uses linear probing to reveal representation drift due to parameter updates. Wu et al. (2021) conducts layer-wise probing on BERT, showing catastrophic forgetting in the top and middle layers. Chen et al. (2023) reveals a correlation between retaining past information and new task learning efficiency through linear probing on k-shot samples. Tao et al. (2023) illustrates BERT’s resilience to catastrophic forgetting even without buffer data. More recently, Zheng et al. (2023b) reveal that most previous work in NLP overlooks the data leakage issue and uses probing techniques to show that LLMs have superior performance on evaluated datasets even before incremental training. Additionally, they reveal that LLMs have strong anti-forgetting ability even under the sequential fine-tuning setting. Our work is inspired by Zheng et al. (2023b) and proposes a novel dataset to minimize the influence of data leakage issues, allowing for correct conclusions about the incremental learning ability of LLMs.
A.2 Understanding Memorization
Fewer studies Carlini et al. (2021); Tirumala et al. (2022); Biderman et al. (2024b); Boix-Adserà et al. (2023) explore memorization within LLMs. For instance, Carlini et al. (2021) discover that GPT-2 can memorize a small proportion of private information during pretraining, raising privacy concerns. Tirumala et al. (2022) show that larger LLMs memorize faster and have higher “forgetting baselines”. Biderman et al. (2024b) reveal the difficulty in predicting which training samples will be memorized by large language models. Boix-Adserà et al. (2023) find that transformers incrementally learn new knowledge, with trained and initial weights progressively increasing in rank. These studies explore memorization from both the perspective of sentences and model weights.
In contrast, this paper studies the problem of memorization and forgetting at a more fine-grained level: the concept level. This approach addresses an underexplored research problem in the incremental learning community.
A.3 Applications in NLP
Many works Sun et al. (2020); Chuang et al. (2020); Liu and Huang (2023); Qiu et al. (2024); Zhang et al. (2023a); Shao et al. (2023a); Zhang et al. (2023b) focus on incremental learning for various NLP tasks, assuming catastrophic forgetting in pre-trained language models and designing techniques to mitigate it. These tasks include text classification Sun et al. (2020); Chuang et al. (2020), relation extraction Liu and Huang (2023), named entity recognition Qiu et al. (2024); Zheng et al. (2022); Zhang et al. (2023a), intent classification Shao et al. (2023a), and machine translation Zhang et al. (2023b).
We refer readers to the survey Zheng et al. (2024) for more applications of incremental learning in NLP tasks. The proposed Concept-1K differs substantially from the aforementioned NLP tasks. It is more challenging and provides better explainability at the concept level.
Appendix B Data Leakage in IL of LLMs
The data leakage issue in NLP is often implicit. Unlike computer vision, where pretraining involves explicit category information, NLP pretraining is self-supervised and lacks clear categorical distinctions that can be easily compared between the pretraining corpus and downstream datasets. This makes it challenging to detect and address data leakage. Therefore, we urge future studies to exercise greater caution regarding data leakage in the IL of LLMs.
B.1 Data Leakage in Classification Tasks
The issue of data leakage in classification tasks for IL with LLMs has recently been highlighted by Zheng et al. (2023b). Their extensive study revisits over 20 IL methods across four key classification tasks: Text Classification, Intent Classification, Relation Extraction, and Named Entity Recognition. One core finding of Zheng et al. (2023b) is that LLMs, such as BERT and GPT-like models, exhibit high probing performance even before they are incrementally trained on specific downstream tasks. This high initial performance suggests that these models already possess substantial knowledge relevant to the classification tasks due to their extensive pre-training on diverse corpora. Consequently, when these LLMs are evaluated under IL settings, the incremental learning of new tasks may, in fact, be leveraging pre-existing knowledge rather than genuinely incremental learning.
This phenomenon leads to misleading conclusions about the effectiveness of various IL methods Sun et al. (2020); Chuang et al. (2020); Liu and Huang (2023); Qiu et al. (2024); Zhang et al. (2023a); Shao et al. (2023a); Zhang et al. (2023b). The high probing performance before task-specific training indicates that the models are not learning incrementally as assumed but rather recalling previously acquired knowledge. Therefore, many IL studies in the context of classification tasks suffer from data leakage, as the benchmark tasks are not truly novel to the LLMs. Addressing this issue requires carefully designed benchmarks that ensure the novelty and exclusivity of the knowledge being tested, a challenge we aim to tackle with our Concept-1K dataset.
B.2 Data Leakage in Generation Tasks
Data leakage poses a significant challenge in IL for generation tasks, which are often more specific and diverse compared to classification tasks. For example, in the IL setting described by Scialom et al. (2022), the task sequence includes Text Simplification, Headline Generation with Constraints, Haiku Generation, Covid QA, Inquisitive Question Generation, Empathetic Dialogue Generation, Explanation Generation, and Twitter Stylometry. Despite the specificity and diversity of these tasks, data leakage remains a concern because LLMs are trained on extensive corpora from the internet and often undergo supervised finetuning on dialogue data OpenAI (2023). This pre-training on vast amounts of internet data means that LLMs might already possess significant knowledge relevant to these generation tasks.
Moreover, the data leakage issue in generation tasks is implicit and easy to overlook. Unlike classification tasks, where techniques like probing Zheng et al. (2023b) can measure LLM performance before training, generation tasks lack such straightforward methods to assess initial model capabilities. This makes it challenging to ascertain how much new knowledge is genuinely being learned during IL versus what the model is simply recalling from its pre-trained knowledge base.
Pioneering work investigating the IL ability of LLMs found that the T0 model Sanh et al. (2021) barely forgets when learning new tasks. This suggests that the T0 model may have already acquired the ability to solve multiple generation tasks with appropriate input prompts before explicit training on them. Detailed results are presented in Figure 2 of Sanh et al. (2021).
Another study by Zhang et al. (2023c) defines generation tasks using different instructions, a paradigm they call continual instruction tuning. They train the T5 model Raffel et al. (2020) sequentially on 19 tasks, achieving 35.7% performance, while jointly training on these tasks yields 42.1%. The close gap between upper bound and lower bound performance indicates minimal forgetting and suggests potential data leakage. Detailed results and experimental settings are provided in Tables 1 and 2 of Zhang et al. (2023c).
Although recent studies Module ; Guo et al. (2024); Huang et al. (2024); Yang et al. (2024); Peng et al. (2024); Ren et al. (2024) claim that forgetting is serious in continual instruction tuning, all of them utilize parameter-efficient finetuning techniques such as LoRA Hu et al. (2021) or prompt tuning Lester et al. (2021). As shown in our experiments in Section 3.3, the IL ability of LoRA and full finetuning differ substantially. LoRA significantly limits the ability to learn new concepts compared to full finetuning, leading to limited new knowledge acquisition and faster forgetting on training samples. Therefore, their IL settings may not accurately reflect the true IL ability of LLMs.
Appendix C Comparison with Existing Incremental Learning Setting
There are three popular IL settings which are widely adopted in the literature of computer vision: class-incremental learning (CIL), task-incremental learning (TIL), and continual pretraining (CPT). However, none of them are appropriate to evaluate the IL ability of LLMs.
C.1 Class-Incremental Learning
CIL is designed for classification tasks such as text classification, and its goal is to learn new classes incrementally. On the one hand, SOTA LLMs with billion-level parameters are overqualified for the above classification tasks with only dozens of categories. On the other hand, the pretraining corpus is likely to contain the knowledge required for the classification tasks (data leakage issue). As shown empirically by Zheng et al. (2023b), sequential training frozen LLMs with expanding classifiers yields comparable or even superior performance with SOTA IL methods.
C.2 Task-Incremental Learning
TIL aims to learn new tasks incrementally Sun et al. (2020); Qin and Joty (2022). Apart from the data leakage issue, the diversity of tasks and orders across research makes it difficult to readily and fairly compare IL algorithms.
C.3 Continual Pretraining
The last scenario CPT aims at continual pretraining models on the corpus from different domains. However, the evaluation relies on the performance of downstream tasks, where we can hardly identify what knowledge is learned or forgotten.
C.4 Summary
In this paper, we consider a novel IL scenario called Instance-level Incremental Learning (IIL). Unlike the IL scenario mentioned above, IIL regards each concept as an instance and is more practical and challenging for existing LLMs. Specifically, we are motivated by the human learning process and expect LLMs to learn new concepts incrementally without forgetting. For example, in Figure 1, humans can learn new concepts that are constantly emerging, such as “Metaverse” and “Quantum Computing”. After learning more concepts such as “Web3.0” and “Non-Fungible Token”, humans will not immediately forget the previously learned concepts such as “Metaverse”.
Appendix D Introduction of Domains in Concept-1K
The domains in Concept-1K are introduced as follows:
D.1 Technology Domain
This domain focuses on both cutting-edge and widely applied technologies. Cutting-edge technologies include artificial intelligence, blockchain, quantum computing, etc., while widely applied technologies encompass cloud computing, the Internet of Things, and more.
D.2 Economic Domain
This domain highlights economic trends and emerging economic concepts. Economic trends cover topics such as digital currency and globalization, whereas emerging concepts include quantitative computing, electronic wallets, peer-to-peer (P2P) networks, and others.
D.3 Education Domain
This domain emphasizes emerging educational technologies and concepts. Technologies such as remote learning and online courses, along with concepts like bilingual education, social education, and lifelong learning, are included.
D.4 Environmental Domain
This domain centers on global environmental issues and green technologies. Topics include climate change, environmental protection, and sustainable energy, as well as technologies like green roofs, shared bicycles, and solar panels.
D.5 Cultural Domain
This domain focuses on diversity and inclusion, and digital media and arts. Diversity and inclusion cover multiculturalism, gender equality, and social inclusion, while digital media and arts include digital art, social media trends, and online communities.
D.6 Health and Medical Domain
This domain is dedicated to emerging medical technologies and public safety. It covers CRISPR gene editing technology, the application of artificial intelligence in medical diagnosis, telemedicine services, wearable health monitoring devices, and concepts related to vaccine development, disease monitoring and prevention strategies, and promoting public health awareness.
Appendix E Concept Selection Criterion
In constructing Concept-1K, we carefully selected concepts based on the following criteria to ensure the relevance, novelty, and richness of the learning material:
E.1 Length criterion
The concept words should not exceed three words in length. This encourages the model to focus on significant and concise terms within the domain, facilitating efficient learning and ensuring that the concepts provide a rich source of information. Shorter concepts are easier to manage and help avoid potential confusion that may arise from overly complex or verbose terms.
E.2 Timeliness criterion
The chosen concept words should preferably be those that emerged after January 2022. This criterion ensures that the general pre-trained models have not yet encountered and learned these concepts and the associated knowledge. By selecting recent concepts, we aim to test the true incremental learning capabilities of LLMs, avoiding biases introduced by prior knowledge.
E.3 Trend criterion
We focus on concepts that are currently receiving widespread attention in academia, industry, and the media. This ensures that the selected concepts are not only relevant and contemporary but also significant and impactful in their respective fields. By choosing trending concepts, we can better gauge the models’ ability to learn and adapt to the latest advancements and discussions in various domains.